Given the function: 6x - 1 2 < 0 63 - f(x) = 62 – 2 x > 0 Calculate the following values: f( - 1) = |-7 f(0) = f(2)
Answer #1

given

f(x)=\left\{\begin{matrix} 6x-1 &\;x<0 \\ 6x-2 &\;x\geq 0 \end{matrix}\right.

we want to calculate the values of f(-1),f(0) and f(2)

1)f(-1)

-1 is less than zero ,so to find f(-1),we use the relation

f(x)=6x-1

(for all values of x less than zero use f(x)=6x-1)

f(-1)=6(-1)-1

f(-1)=-6-1

f(-1) = -7

2) f(0)

to find this use the second relation, because for the values of x greater than or equal to zero,use second relation.

f(x)=6x-2

f(0)=6*0-2

f(0)=0-2

f(0)=-2

3) f(2)

to find this use the second relation, because for the values of x greater than or equal to zero,use second relation.

f(x)=6x-2

f(2)=6*2-2

f(2)=12-2

f(2)=10